OpenStudy (anonymous):

@dan815 Could you check my answer?

OpenStudy (anonymous):

The volume of a cube is 42 cubic feet. What is the volume of a pyramid that fits exactly inside the cube?

OpenStudy (anonymous):

What is it

OpenStudy (anonymous):

In order to figure out the volume of the pyramid that fits exactly inside the cube, we must first examine the formula's for both the cube and the pyramid. Volume of a Cube: V = (b) (h) Volume of a pyramid: V = 1/3 (b) (h) As you can see, the volume of the pyramid is 1/3 the volume of the cube, again, like the cylinder and the cone, the volume of the pyramid is 1/3 the volume of the cube, so in order to find the volume we must divide the volume of the cube by three, lets do that. 42 / 3 = 14 As you can see, 42 / 3 equals 14, which means that the volume of the pyramid is also 14.

OpenStudy (anonymous):

You are correct.

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

The volume of a cube is actually a^2, but it's the same as (b) (h)

OpenStudy (anonymous):

We just have to multiply 1/3 or divide by the reciprocal 3.